# Knowing Your Nonlinearities: Shapley Interactions Reveal the Underlying Structure of Data

### v1

@@ Abstract
Measuring nonlinear feature interaction is an established approach to understanding complex patterns of attribution in many models. In this paper, we use Shapley Taylor interaction indices (STII) to analyze the impact of underlying data structure on model representations in a variety of modalities, tasks, and architectures. Considering linguistic structure in masked and auto-regressive language models (MLMs and ALMs), we find that STII increases within idiomatic expressions and that MLMs scale STII with syntactic distance, relying more on syntax in their nonlinear structure than ALMs do. Our speech model findings reflect the phonetic principal that the openness of the oral cavity determines how much a phoneme varies based on its context. Finally, we study image classifiers and illustrate that feature interactions intuitively reflect object boundaries. Our wide range of results illustrates the benefits of interdisciplinary work and domain expertise in interpretability research.
@@ 1 Introduction
Feature attribution is a common approach to interpreting modern neural networks. One classic method is the Shapley decomposition (Shapley, 1952), which is adapted from game theory scenarios where the goal is to attribute credit—a Shapley value—within the context of a multi-agent cooperative game. The Shapley decomposition assumes approximately linear features, usually an incorrect assumption where deep learning is concerned. Consequently, researchers have often tried to quantify the validity of the approximate linearity assumption in Shapley values (Kumar et al., 2021). These methods calculate the degree of nonlinearity, or Shapley interaction, among a set of features; itself a metric that can be used to interpret the representations produced by a complex model.
This paper investigates Shapley interactions in a number of tasks and architectures. We use Shapley interactions as a case study to illustrate the importance of grounding model interpretations in the underlying structure of the data and the target models. To this end, we draw connections between interaction metrics and various structural properties of the data in each setting: syntax, tokenization, and idiomatic expressions in masked and autoregressive language models (MLMs and ALMs, respectively); phoneme articulation differences in speech models; and distinctions between edges, foreground, and background pixels in image classifiers. After introducing our approach to Shapley interactions, we apply them in a variety of settings and find the following.
  • When we control for positional distance (Section 3.1), Transformer-based MLMs—but not ALMs—show a strong correlation between feature interaction and syntactic proximity on a pair of context tokens (Section 3.2). Both LMs often exhibit stronger interactions between pairs of context tokens within an idiomatic Multiword Expression (MWE), but the pattern (Section 3.3) is more consistent in MLMs predicting nearby tokens and in ALMs predicting distant tokens. This combination of observations would suggest that the structure of nonlinear interactions learned by MLMs is more overtly hierarchical than that of ALMs.
  • It is known to phonologists that the acoustics of a vowel cannot be interpreted in isolation because the vocal tract is shaped by nearby consonants (Rakerd, 1984). We find that acoustic features in speech models accordingly interact more around transitions between consonants and vowels compared to transitions between two consonants (Section 4.1). Furthermore, consonance have more nonlinear interactions on average between consecutive acoustic features near their transition if the consonant is articulated with a more open oral cavity similar to a vowel, rather than a closed oral cavity (Section 4.2).
  • In image classifiers, pixels close to object boundaries exhibit less local interaction, likely because any perturbations are obscured by nearby edge compression artifacts (Section 5.1). Considering the edge pixels themselves in detail, edges interact more with nearby foreground object pixels than with other nearby pixels, but interact similarly with all distant pixels (Section 5.2). We may infer that the boundary of the object is determined by both foreground and edge pixels, but at a distance other interactions take precedence, such as the edge-edge interactions that determine the shape of an edge overall.
2 Background: Shapley Interactions
Shapley values are used to attribute decisions to specific features in predictive models, ideally by exhaustively evaluating each possible coalition of interacting features. Specifically, the Shapley value of a feature is obtained by computing the difference in a model’s output when a feature is included versus when it is withheld from a given set.
Formally, Shapley values are calculated by taking the marginal contribution of a set of target features $A$ to each subset $S\subseteq N\backslash A$, where $N$ denotes the set of all features. For a value function $v$, here the logit outputs of a neural network, the Shapley value is the weighted average across all possible subsets when ablating $A$, given by:
$\phi(A)=\sum_{S\subseteq N\backslash A}\frac{|S|!(|N|-|S|-1)!}{|N|!}(v(S\cup A)-v(S))$ (1)
Shapley values decompose into a close approximation of the output when interactions are additive, such that $\phi(\emptyset)\approx\sum_{i\in S}v(\{i\})$. In scenarios where features are dependent and their composition is non-linear, Shapley values do not account for interacting effects between coalitions. The violation of Shapley assumptions can be computed through the Shapley residual (Kumar et al., 2021):
$r_{i}=\nabla_{i}\phi-\nabla\phi(\{i\})$ (2)
However, calculating the Shapley values for each coalition requires iterating over the powerset of $N$, requiring $2^{d}$ calculations, where $d$ is the total number of features. In high-dimensional input spaces such as the output vocabulary of a language model, the exact calculation of Shapley residuals is therefore prohibitively expensive. Instead, we must rely on a first-order approximation of the interactions based on the Shapley Taylor interaction index (STII) (Agarwal et al., 2019) using the discrete second-order derivative. For simplicity, we consider the case of interaction between a pair of feature sets $A$ and $B$. Similar to Saphra & Lopez (2020), we scale the residual by the norm of the entire sequence with no feature ablations.
$\textrm{STII}_{A,B}=\frac{\|\phi(\emptyset)-\phi(A)-\phi(B)+\phi(A,B)\|_{2}}{\|\phi(\emptyset)\|_{2}}$ (3)
Most often, Shapley values are scalar values, but the influence on each unit in a vector can be computed individually to form a vector of Shapley values. Taking this approach to compute a vector of interactions, we then use norm to ultimately convert it to back a scalar metric of interaction.
3 Language Modeling
@@
Our first experiments are on language models, measuring how known associations between tokens correlate with Shapley-based measures of feature interaction. We consider the influence of token position, idiomatic phrases, and syntax. We find that MLMs and ALMs differ in their interaction structure, especially in how they respond to syntax.
Datasets
@@
All English language modeling experiments use wikitext-2-raw-v1 (Merity et al., 2016) tokenized and dependency parsed (for syntax experiments) with spaCy (Honnibal et al., 2020). For MWE experiments, we use the AMALGrAM (Schneider et al., 2014a) supersense tagger, which identifies both strong and weak (Schneider et al., 2014b) MWEs. For syntax experiments, we resolve incompatibilities between the spaCy tokenizer and the model-specific tokenizers by assigning overlapping tokens a syntactic distance of zero.
Experimental Setup
We analyze two models, the ALM GPT-2 (Radford et al., 2019) and the MLM BERT-base-uncased (Devlin et al., 2018). We apply softmax to logit outputs to ensure interactions across examples are comparable. Each sample sentence is unpadded and truncated to 50 tokens. The first and second tokens in an interacting pair are, respectively, $x_{t_{1}}$ and $x_{t_{2}}$ where $t_{1}$ and $t_{2}$ provide the indices of the features in an input sequence. We measure their non-linear interactions by the Shapley interaction of two feature sets containing a single token each, $\textrm{STII}_{\{x_{t_{1}}\},\{x_{t_{2}}\}}$. We denote the index of the target token to be predicted as $t_{target}$. We ablate the tokens by replacing them with the padding token of the tokenizer.
Figure 1: Average interaction scores monotonically decrease with greater interacting token pair distance $d_{i}$ and with a greater prediction distance $d_{p}$.
3.1 Positional distances
One potential factor influencing interactions between tokens is the positional distance between the interacting pair, where it is likely that stronger interactions occur between closer tokens. We also consider the positional distance between the interacting pair and the target of prediction. These surface-level positional differences are the dominant factor in feature interaction.
When relating STII to the latent structure of a sequence, we control for both interacting pair distance $d_{i}$, i.e., positional distance between the interacting pair (Equation 4); and prediction distance $d_{p}$, i.e., the positional distance between the pair and the predicted target token (Equation 5).
@@ 3.1 Positional distances
3.1.1 Results
@@
We confirm from Figure 1 that in both ALMs and MLMs, STII monotonically decreases at greater prediction and interacting pair distances. The dramatic decline of STII with increased prediction distance implies that when these models predict tokens, they treat the more distant context as a bag of words rather than as complex syntactic relations (Khandelwal et al., 2018). We also see that closer tokens interact more strongly with each other. In general, interactions for small distances are much higher for the autoregressive model, suggesting that the MLM behaves in a more linear and strictly compositional manner. Our other experiments stratify samples by these distances when we study other factors influencing feature interactions, as it is clear that higher interacting pair distance and positional distance both reduce the interaction score in MLMs and ALMs.
(a) ALM experiments
(b) MLM experiments
Figure 2: Each cell represents the Spearman correlation between syntactic distance and STII, for a given interacting pair distance and prediction distance. Each syntactic distance included must have at least 50 data points. We only provide results for cells where there exists at least one direct syntactic modifier pair separated by the positional distance $d_{i}$ and the Spearman correlation given at that cell is statistically significance ($p<0.05$).
3.2 Syntactic distance
@@
Syntactic structure can also influence an LM’s predictions. If a model composed distant syntactic relations in a linear way, it would treat the wider context as though it were a bag of words. By instead exhibiting strong interactions between syntactically close tokens, the model would closely entangle the meaning of a modifier with its head. We measure syntactic distance by the number of dependency edges traversed to connect a pair of tokens, a metric encoded by projected representations in both MLMs (Hewitt & Manning, 2019) and ALMs (Murty et al., 2022). We verify the role of modifier connections by the Spearman correlation between syntactic distance and STII, stratified by interacting pair distance and prediction distance. In line with previous findings on LSTM ALMs (Saphra & Lopez, 2020), we expect the STII to negatively correlate with syntactic distance.
3.2.1 Results
Figure 2 shows correlation between syntactic distance and STII. We confirm that all statistically significant correlations are negative. These correlations are stronger in MLMs, particularly for distant interacting pairs, suggesting that the latent syntactic structure is limiting the model’s tendency to reduce interactions between distant tokens. This finding suggests we may view feature interaction as an overtly syntactic alternative to the native encoding of distance that occurs in autoregressive architectures (Haviv et al., 2022). The degree of interaction is learned to be more syntactic, rather than relying on positional information in the language modeling objective itself.
@@ 3.3 Multiword Expressions
(a) ALM experiments
(b) MLM experiments
Figure 3: Each plot represents the trend in STII for a given prediction distance $d_{p}$, responding to the x-axis of interacting pair distance $d_{i}$. Trend is given for the average over: all pairs, strong MWE relations, and weak MWE relations.
Classical treatments of semantics are compositional, implying that the meaning of a sentence is derived by composing the meanings of each individual word. However, there are groups of words whose meaning can only be derived when looking at the entire group rather than the individual words. These word groups, known as multiword expressions (MWEs), include idioms like break a leg, where the isolated meaning of each of the component words break, a, and leg fail to compose the meaning of the entire expression. Higher interaction values for the tokens in the idiom would indicate a less compositional treatment of the whole phrase.
@@
3.3.1 Results

### v2

# Using Shapley interactions to understand how models use structure

@@ Abstract
Language is an intricately structured system, and a key goal of NLP interpretability is to provide methodological insights for understanding how language models represent this structure internally. In this paper, we use Shapley Taylor interaction indices (STII) in order to examine how language and speech models internally relate and structure their inputs. Pairwise Shapley interactions measure how much two inputs work together to influence model outputs beyond if we linearly added their independent influences, providing a view into how models encode structural interactions between inputs. We relate the interaction patterns in models to three underlying linguistic structures: syntactic structure, non-compositional semantics, and phonetic coarticulation. We find that autoregressive text models encode interactions that correlate with the syntactic proximity of inputs, and that both autoregressive and masked models encode nonlinear interactions in idiomatic phrases with non-compositional semantics. Our speech results show that inputs are more entangled for pairs where a neighboring consonant is likely to influence a vowel or approximant, showing that models encode the phonetic interaction needed for extracting discrete phonemic representations.
Using Shapley interactions to understand how models use structure
       Divyansh Singhvi11footnotemark: 1        Independent        divyanshsinghvi@gmail.com                                   Diganta Misrathanks: equal contribution            ELLIS & MPI-IS Tübingen            diganta.misra@tue.ellis.eu                                      Andrej Erkelens11footnotemark: 1               Independent               andrej.erkelens@gmail.com                        Raghav Jain11footnotemark: 1 UC San Diego r6jain@ucsd.edu
Isabel Papadimitriouthanks: equal contribution Kempner Institute, Harvard University isabelpapadimitriou@fas.harvard.edu                        Naomi Saphra22footnotemark: 2 Kempner Institute, Harvard University nsaphra@fas.harvard.edu
@@ 1 Introduction
How do language model features work together to influence prediction results? Do the internals of language models reflect the complex structure of language in how they combine features? Feature attribution—measuring how different model features (like inputs or neurons) influence output decisions in isolation—is a key method of understanding and interpreting neural models. One common approach to feature attribution is adapted from game theory scenarios, and treats features like agents in a cooperative game, attributing credit for the outcome to each feature (Lundberg and Lee, 2017). This credit value, or Shapley value (Shapley, 1952), quantifies the additive effect of each feature on the output, assuming that features act in a linearly independent manner on the output. The linearity assumption is not accurate for most deep learning scenarios: neural networks are non-linear, and features interact in complex ways inside model representations to influence output predictions.
What interactions between features do we miss when we assume this linear independence? To address this question, researchers have proposed methods to calculate how much information we lose when assuming additivity (Kumar et al., 2021), and Shapley interactions, accounting for how features have influence in pairs or groups on top of how they act independently (Agarwal et al., 2019).
In this paper, we investigate how Shapley interactions can enhance our understanding of the internal processes of language models. We ground our investigation in structural features that we know about the input data (like syntactic structure), and ask: what do Shapley interactions reveal about how the model uses the dependency structure in language? By relating Shapley interactions to structural linguistic features, we showcase how different models use (or don’t use) linguistic structural features in their internal representations. We run experiments on autoregressive and masked text models, as well as on automatic speech recognition models, and report the following findings:
  • Autoregressive models (but not masked models) show a strong correlation between Shapley interaction and the syntactic proximity of features. This result indicates that syntactic structure is encoded in non-linear interactions between model features (Section 3.2).
  • Both autoregressive and masked models exhibit stronger interactions between pairs of tokens in multiword expressions (MWEs) that have idiomatic non-compositional meaning (expressions like to kick the bucket, meaning to die) (Section 3.3).
  • In speech models, Shapley interactions are stronger between consonants and vowels than between pairs of consonants (Section 4.1). This result accords with the known phenomenon of coarticulation: the acoustics of vowels are often shaped by the surrounding consonants, while consonants can be understood in isolation (Rakerd, 1984). This finding also extends to more sonorant vowel-like consonants, which interact more with surrounding consonants than those produced with the vocal tract more closed (Section 4.2).
The non-linearities and interactions in model internals are a vital missing piece of the wider language model interpretability inquiry. Our work showcases Shapley interactions as a powerful interpretability method for examining how language models organize their representations to reflect latent input structure.
2 Background and related work
Figure 1: The effect of interacting pair distance and prediction distance on feature interactions. Our results for the experiments relating Shapley Interactions with a token’s position in the sequence. We find that, for both autoregressive models (left) and masked models (right), STII decreases monotonically with distance. This holds both when we are measuring distance as the distance between the two elements in the interacting pair ($d_{i}$, blue line) and when we are measuring distance between the interacting pair and the token that the model is predicting ($d_{p}$, orange line). Our results indicate that models treat tokens that are far away from each other more like an unentangled bag-of-words, and that they treat pairs of tokens that are far away from the token being predicted as unentangled, no matter the distance between them.
2.1 Shapley Interactions
Shapley values are used to attribute decisions to specific features in predictive models. The Shapley value of a set of features $A$ is obtained by computing the difference in a model’s output when $A$ is included versus when it is excluded. If we take the set of all features $N$ and remove $A$, we want to see how much value $A$ adds to every possible context subset $S\subseteq N\backslash A$. In our case, the value function $v$ is the logit output of the model. The Shapley value is the weighted average of this marginal contribution over all $S$:
$\begin{split}&\phi(A)=\sum_{S\subseteq N\backslash A}w_{S}\left(v(S\cup A)-v(S)\right)\end{split}$ (1)
where the weight $w_{S}$ for each subset is the number of possible subsets $S$ of the same size:
$w_{S}={{|N|-|A|}\choose|S|}$ (2)
If the features are linearly additive without interactions, note that $\phi(\emptyset)\approx\sum_{i\in N}v(\{i\})$. However, in scenarios where features are dependent and their composition is non-linear, Shapley values do not account for interacting effects between sets. Various methods to quantify and use these interactions have been proposed (Owen, 1972; Grabisch and Roubens, 1999; Fumagalli et al., 2023; Tsai et al., 2023; Kumar et al., 2021).
For simplicity, we consider only pairwise interactions between two feature sets $A$ and $B$. To calculate pairwise Shapley interactions, we rely on the Shapley Taylor interaction index (STII) (Agarwal et al., 2019) to calculate second-order interactions using the discrete second-order derivative. Since our features are vectors, we calculate the scalar Shapley interaction value for each dimension individually, and use the norm of this vector as an interaction metric. Similar to Saphra and Lopez (2020), we scale the result by the norm of the entire sequence with no feature ablations.
$\textrm{STII}_{A,B}=\frac{\lVert\phi(\emptyset)-\phi(A)-\phi(B)+\phi(A,B)\rVert_{2}}{\lVert\phi(\emptyset)\rVert_{2}}$ (3)
Calculating the Shapley values for each coalition requires iterating over the powerset of $N$, requiring $O(2^{|N|})$ calculations. In high-dimensional input spaces, the exact calculation of Shapley residuals is therefore prohibitively expensive. We compute the Shapley value using the method of Mitchell et al. (2022), which approximates the weighted average of subset credits with Monte Carlo Permutation Sampling (Castro et al., 2009).
2.2 Structure in language models
There is a huge and varied literature aimed at understanding how language models use and represent the structure in their input. Approaches include examining if the output probabilities of language models reflect structural rules (Warstadt et al., 2018; Hu et al., 2024; Gauthier et al., 2020, inter alia), as well as looking inside model representations. For the latter approaches, while many linguistic structural elements can be linearly extracted from the representations of text and speech models (Hewitt and Manning, 2019; Belinkov, 2021; Pasad et al., 2024; Chrupała et al., 2020; Park et al., 2023, inter alia), and attribution methods can relate the linear importance of different features in text and speech models (Markert et al., 2021; Ethayarajh and Jurafsky, 2021; Yeh et al., 2020; Kokalj et al., 2021, inter alia), the fact remains that neural models have complex nonlinearities in their internal processing.
How can we analyze the ways in which nonlinear interactions play out in model internals, and what they encode? Multiple papers have analyzed the difficulties of knowing what we can extract when using nonlinear probing methods (Voita and Titov, 2020; Pimentel and Cotterell, 2021; Hewitt et al., 2021), and others have proposed searching for causal effects which can be generally agnostic to whether the processing is linear (Geiger et al., 2021; Arora et al., 2024). Shapley interactions let us directly link features of the input to different extents of nonlinear processing. Prior work showing the utility of Shapley interactions in analyzing NLP models has focused on older architectures like LSTMs (Saphra and Lopez, 2020; Singh et al., 2019) and on models fine-tuned for simple text classification tasks (Jumelet and Zuidema, 2023; Chen et al., 2020). Our work builds on and generalizes these results by relating Shapley interactions to diverse forms of linguistic structure (syntactic, semantic, and phonetic) on models trained on domain-general language tasks (generation for text, and ASR for speech)
3 Text models: Interactions between tokens
@@
Our first experiments are on language models, measuring how known associations between tokens correlate with Shapley-based measures of feature interaction. We consider the influence of token position, idiomatic phrases, and syntax. We find that masked LMs and Autoregressive LMs differ in their interaction structure, especially in how they respond to syntax.
Models and Datasets
We run all of our experiments on two models: the autoregressive model GPT-2 (Radford et al., 2019) and the masked language model BERT-base-uncased (Devlin et al., 2018). Each input sentence is unpadded and truncated to 20 tokens, and we apply softmax to the logit outputs to ensure that interactions across different examples are comparable.
@@
All English language modeling experiments use wikitext-2-raw-v1 (Merity et al., 2016) tokenized and dependency parsed (for syntax experiments) with spaCy (Honnibal et al., 2020). We resolve incompatibilities between the spaCy tokenizer and the model-specific tokenizers by assigning overlapping tokens a syntactic distance of zero. For the multiword expression experiments, we use the AMALGrAM supersense tagger (Schneider et al., 2014a), which identifies both strong and weak (Schneider et al., 2014b) MWEs.
(a) Autoregressive language model experiments
(b) Masked language model experiments
Figure 2: Syntactically close token pairs interact more than other pairs with the same positional distance. The results of our syntactic distance experiments (Section 3.2): how does syntactic distance correlate with STII, controlling for the effect of position? A negative correlation means that tokens closer in the parse tree (low syntactic distance) are more heavily entangled (high STII). Autoregressive models show a consistently negative correlation in all significant cells, meaning that syntax is encoded in Shapley interactions. We stratify our results by the two positional distance metrics in Section 3.1, so that we can calculate the effect of syntactic distance, marginalizing out the effect of positional distance. Each cell displays a correlation between syntactic distance and STII for a given interacting pair distance and prediction distance. We only provide results for cells where there exists at least one direct syntactic modifier pair separated by the positional distance $d_{i}$ and the Spearman correlation given at that cell is statistically significance ($p<0.05$). For our correlation calculation, we only include a syntactic distance if there are at least 50 data points with that syntactic distance in our data set.
3.1 Baseline: the effect of position
One potential factor influencing interactions between tokens is the positional distance between tokens Let’s say that we are calculating the interaction between two tokens, $x_{t_{1}}$ and $x_{t_{2}}$ at positions $t_{1}$ and $t_{2}$. The token that the model is trying to predict (i.e. the next token in autoregressive models, and the masked token in masked models) is at position $t_{\text{target}}$. There are two relevant positional distances that are likely to influence interaction.
Firstly, the interacting pair distance, $d_{i}$, is the distance between the two tokens:
@@ 3.1 Baseline: the effect of position
Secondly, the prediction distance, $d_{p}$, is the distance between the pair of tokens that we are calculating the interaction of, and the target token that the model is trying to predict:
@@
For our position baseline experiments, we test how both interacting pair distance and prediction distance influence the STII between the two tokens $x_{t_{1}}$ and $x_{t_{2}}$.
Results
@@
Our results are presented in Figure 1, confirming that distance has an effect on STII in both autoregressive and masked models. This holds whether we are measuring distance as distance between the interacting pair (interacting pair distance $d_{i}$) or distance between the last token in that pair and the target prediction token (prediction distance $d_{p}$). The dramatic decline of STII with increased prediction distance implies that when these models predict tokens, they treat the more distant context as a bag of words rather than as complex syntactic relations (Khandelwal et al., 2018). We also see that closer tokens interact more strongly with each other.
For the rest of our experiments, we will stratify samples by both $d_{i}$ and $d_{p}$, so that we can measure the effects of linguistic structure beyond these position effects that we demonstrate here.
(a) Autoregressive LM experiments
(b) Masked LM experiments
Figure 3: Tokens interact more if they occur in the same multiword expression: Shapley interactions are higher for tokens in multiword expressions than tokens that are not. The results are controlled for prediction distance $d_{p}$ (different facets) and interacting pair distance $d_{i}$ (x-axis). Within each facet for each x-axis value, we can see that the STIIs for tokens in Strong MWEs (blue) and Weak MWEs (orange) are significantly higher than the average over all pairs (green).
3.2 Syntactic structure
@@
Syntactic structure can also influence an LM’s predictions. If a model composed distant syntactic relations in a linear way, it would treat the wider context as though it were a bag of words. By instead exhibiting strong interactions between syntactically close tokens, the model would closely entangle the meaning of a modifier with its head. We measure syntactic distance by the number of dependency edges traversed to connect a pair of tokens, a metric encoded by projected representations in both masked (Hewitt and Manning, 2019) and autoregressive (Murty et al., 2022) models. We verify the role of modifier connections by the Spearman correlation between syntactic distance and STII, stratified by interacting pair distance and prediction distance.
Results
Figure 2 shows correlation between syntactic distance and STII. Our analysis reveals that, for autoregressive language models, all statistically significant correlations are negative. In contrast, non-autoregressive language models exhibit both positive and negative correlations. This finding aligns with Saphra and Lopez (2020)’s research on LSTMs showing that syntax is handled more consistently in autoregressive models, and with Ahuja et al. (2024), who in a different setting show that autoregressive models are more predisposed to syntax-style generalizations.
The inconsistencies observed in non-autoregressive models may stem from their handling of positional proximity in less intuitive ways, complicating the relationship between syntactic and linear distance. The interaction between these two dimensions may be more difficult to manage in masked models, leading to the varied correlation outcomes.
This finding suggests that we can interpret feature interaction as a distinctly syntactic alternative to the inherent distance encoding found in autoregressive architectures (Haviv et al., 2022). In these models, the degree of interaction is learned to prioritize syntactic relationships rather than depending solely on positional information within the language modeling objective. This highlights a fundamental difference in how these models integrate syntactic structure and distance.
@@ 3.3 Multiword expressions
While semantics is often treated as compositional (the meaning of a sentence can be composed by rules, following the syntax and the meaning of each individual word), language is also characterized by non-compositional, or idiomatic, phrases. These are groups of words whose meaning can only be derived when looking at the entire group rather than the individual words. These word groups, known as multiword expressions (MWEs), include idioms like break a leg, where the isolated meaning of each of the component words break, a, and leg fail to compose the meaning of the entire expression. Higher interaction values for the tokens in the idiom would indicate a less compositional treatment of the whole phrase.
We use the MWE tagger from Schneider et al. (2014a), which also distinguishes strong and weak MWEs. As explained by Schneider et al. (2014c), strong MWEs are non-compositional idiomatic phrases, in which the meaning of each word is changed substantially by its context. Weak MWEs, meanwhile, are sequences with distributional properties similar to strong MWEs, but where the meaning of each word is maintained. In other words, near miss would be a weak MWE—it means an event that was nearly a miss—but close call would be a strong MWE, as the meanings of close and of call are both tied to the particular idiom.
@@
Figure 4: Vowel-consonant interactions are higher than consonant-consonant interactions. Average STII between pairs of adjacent acoustic input features, adjusting the interval range around the phone boundary. Confidence intervals are provided by bootstrap.
Results